Given the following information for a portfoilio of 60 % in stock and 40% in bonds calculate:

1) The expected return of the portfolio

2) The risk of the portfolio Scenario

Scenario

Prob.

Stock Fund

ROR

Bond Fund

ROR

Severe Recession

0.10

-25

-10

Mild Recession

0.20

-8

1

Normal Growth

0.45

17

5

Boom

0.25

34

8
Answer #1

1). Expected Return = =1 [Probability(i) * Return(i)]

Stock's Expected Return = [0.10 * -25%] + [0.20 * -8%] + [0.45 * 17%] + [0.25 * 34%]

= -2.50% - 1.60% + 7.65% + 8.50% = 12.05%

Bond's Expected Return = [0.10 * -10%] + [0.20 * 1%] + [0.45 * 5%] + [0.25 * 8%]

= -1.00% + 0.20% + 2.25% + 2.00% = 3.45%

Portfolio's Expected Return = [0.60 * 12.05%] + [0.40 * 3.45%] = 7.23% + 1.38% = 8.61%

2). Portfolio's Standard Deviation = [=1{Probability(i) * (Expected Return - Return(i))}2]1/2

= [{0.6 * (8.61% - 12.05%)}2 + {0.4 * (8.61% - 3.45%)}2]1/2

= [7.1002%2 + 10.6502%2]1/2 = [17.7504%2]1/2 = 4.21%